OpenStudy (nicole143):

What is the sum of the geometric series? I am very confused on how to do this. If anyone could help me understand (Go through the steps), that would be great! I'll put the equation put in one second..

OpenStudy (nicole143):

\[\sum_{n-1}^{10} 6(2)^2\]

OpenStudy (nicole143):

Oops \[\sum_{n-1}^{10} 6(2)^n\]

OpenStudy (mathmale):

First of all, here's a reference that you might find useful: http://www.purplemath.com/modules/series5.htm There are 2 cases here: 1) find the sum of the first n terms of the series, and 2) find the sum of the infinite series (n goes to infinity). In your case you're going from n=1 to n=10. I ask you to let n=1 right off. What is the value of the first term of this series?

OpenStudy (mathmale):

Nicole: are you sure that you mean n-1? Could you possibly have meant n=1?

OpenStudy (nicole143):

Yes, n=1 Sorry ha - Thought I fixed all the mistakes

OpenStudy (mathmale):

:)

OpenStudy (mathmale):

So, if n=1, what is the first term of the series?

OpenStudy (nicole143):

I don't even know what \[\sum_{ }^{ }\] means so if you could first tell me that, it would be very helpful :P

OpenStudy (mathmale):

That's the greek letter, sigma. In Math, it's a command; it tells you to add up everything that follows it. If you write SIGMA, starting at n=1 and ending at n=5, of n, you get \[\sum_{n=1}^{5}n=1+2+3+4+5=15.\]

OpenStudy (nicole143):

Oh, I see. Okay so the first term would be 1 and then +2 + 3 + 4 + 5 + 6 + 7 + 8 + 9 + 10 ??

OpenStudy (mathmale):

The problem you've posted is \[\sum_{n=1}^{10}6(2)^n\] which is a bit easier to do if you take that factor 6 outside: \[6\sum_{n=1}^{10}2^n\]

OpenStudy (mathmale):

Yes, to your question. Add up all 10 of those n values, and then multiply the resulting sum by 6. Done.

OpenStudy (nicole143):

55 * 6 = 330 ?

OpenStudy (mathmale):

Yes, that's great. Right on! But this is for your EXAMPLE. The problem you've posted is a bit different: \[6\sum_{n=1}^{10}2^n\] Here we still let n=1, 2, 3, ...., 9, 10, producing 2^1, 2^2, ... So y ou will obtain 10 powers of 2. You could calculate those and add them up, but boy, would that be slow!

OpenStudy (nicole143):

Okay, so 330 is not eh finial answer. Good, I was confused ha

OpenStudy (mathmale):

http://www.purplemath.com/modules/series5.htm provides you with a formula for the sum of a geometric series. It looks like: \[\sum_{i=1}^{n}r^n=\frac{ 1-r^n }{ 1-r }\] Can y ou find that formula in this web page?

OpenStudy (nicole143):

Yes

OpenStudy (mathmale):

Good. Now, are you familiar with the term "common ratio"? In the posted problem, your common ratio is 2. Can you explain why that is?

OpenStudy (nicole143):

Because that is were "r" would be

OpenStudy (nicole143):

*where

OpenStudy (nicole143):

I did the adding thing just to see but I didn't get the right answer :( @mathmale

OpenStudy (mathmale):

The problem you posted is \[\sum_{i=1}^{10}6(2)^n\] and the common ratio is r. This means that every term equals the preceding term, times 2.

OpenStudy (mathmale):

Please go back and copy onto paper the formula I gave you for the nth partial sum. Let me know when you've done that.

OpenStudy (nicole143):

\[\sum_{i=1}^{n} r^2 = \frac{ 1 - r^2 }{ 1 -r }\] this one?

OpenStudy (mathmale):

Yes, except that it's r^n, not r^2. Are you clear on that? We will substitute r=2 later.

OpenStudy (nicole143):

Oh, okay. So do I now fill in? Or no?

OpenStudy (mathmale):

Now, replace that r with " 2 " Replace that n with 10 (on the right side ONLY). What expression do you get ? Don't try to evaluate it.

OpenStudy (nicole143):

\[\sum_{i-1}^{n}r^2 = \frac{ 1-2^{10} }{ 1-2 }\]

OpenStudy (mathmale):

that's right. now, because you also had a ' 6 ' in the problem you posted, multiply both sides of your equation by 6. Note that the left side should look like\[6\sum_{i=1}^{10}2^n\]

OpenStudy (mathmale):

where r is the common ratio and 10 is the number of terms you're adding up.

OpenStudy (mathmale):

What do you get on the right side now?

OpenStudy (nicole143):

Okay, um 1,023 ??

OpenStudy (nicole143):

Then times 6 to have a final answer of 6,138 ?? :D

OpenStudy (mathmale):

I'd just like to see:\[6\frac{ 1-2^{10} }{ 1-2 }\]

OpenStudy (mathmale):

By reversing the order of subtraction, you could also write\[6\frac{ 2^{10}-1 }{ 2-1 }\]

OpenStudy (nicole143):

Okay, so is 6,138 wrong then?

OpenStudy (mathmale):

which you could either simplify or leave as is.

OpenStudy (mathmale):

Note that \[\frac{ 6 }{ 2-1 }=6.\]

OpenStudy (mathmale):

What does this leave you with?

OpenStudy (nicole143):

6,138 ??

OpenStudy (mathmale):

Looks OK to me (I haven't actually done the calculation). How did you evaluate 2^10?

OpenStudy (nicole143):

2 to the power of 10 on my calculator

OpenStudy (mathmale):

Great. You've done very well here. I'd suggest you review what we've discussed and be sure you can do this problem unaided. Although i'd like to get off the 'Net now, I'd welcome further questions. Others are capable of answering them too, if you want fast results; I may be back on later this evening or sometime tomorrow morning.

OpenStudy (nicole143):

Have a good night! Thank you!

OpenStudy (mathmale):

My pleasure. Take care, Nicole!